A food manufacturer wishes to estimate the average time required to process orders

received from supermarkets. A random sample of 60 orders revealed a mean time of 2.4

days and a standard deviation of 1.2 days.

i) Compute a 95% confidence interval for the population mean process time.

ii) If the manufacturer wishes to be within 0.2 days of the true average time, what size

sample should be drawn?



1
Expert's answer
2021-08-02T16:03:14-0400

i) The critical value for "\alpha=0.05" and "df=n-1=60-1=59" degrees of freedom is "t_c= 2.000995." The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(2.4-2.000995\times\dfrac{1.2}{\sqrt{60}}, 2.4+2.000995\times\dfrac{1.2}{\sqrt{60}})"

"=(2.09, 2.71)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "2.09<\mu< 2.71," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(2.09, 2.71)."


ii)


"t_c\times\dfrac{s}{\sqrt{n}}\leq0.2"

"n\geq(\dfrac{t_c\times s}{0.2})^2"

"n\geq(\dfrac{t_c\times 1.2}{0.2})^2"

"n\geq 36(t_c)^2"

"n=141, df=140, t_c=1.977054"



"36(1.977054)^2=140.7"

"n=140, df=139, t_c=1.977178"



"36(1.977178)^2=140.7"

Hence

"n=141"